I would like to know how to factor 64x^3-27y^3. Also, i need to be provided with the steps to getting to my solution as well! Thank you
Answer by AnlytcPhil(1808)

We memorize the way to factor the sum or the difference of cubes:

A³ ± B³ = (A ± B)(A² ∓ AB + B²)

That is, the sum of two cubes factors like this:

A³ + B³ = (A + B)(A² - AB + B²)

And the difference of two cubes factors like this:

A³ - B³ = (A - B)(A² + AB + B²)

Be sure to memorize these. 

Your problem is

64x³ - 27y³

since 64 = 4³ and 27 = 3³, we can write that as

(4x)³ - (3y)³

So we substitute A = 4x  and B = 3y  in

A³ - B³ = (A - B)[A² + AB + B²]

(4x)³ - (3y)³ = (4x - 3y)[(4x)² + (4x)(3y) + (3y)²]

Then we simplify the bracket expression:

              = (4x - 3y)[16x² + 12xy + 9y²]
 
              = (4x - 3y)(16x² + 12xy + 9y²)

Edwin